suppose a monopolist sells to two groups that have constant elasticity demand curve with elasticity e1 and e2. the marginal cost of production is constant at C. What price is charged to each group?


1
Expert's answer
2021-06-01T11:27:17-0400

If the monopoly firm is selling its product to two different customer groups and if it is charging a different price to each of the customer groups for the same or identical good then the monopoly firm is said to be engaged in the practice of third-degree price discrimination.

The price elasticity of demand measures the proportionate change in the quantity demanded of a good due to some proportionate change in the price of the good keeping constant all other factors that are able to affect the demand of the good.

As monopoly firm knows that the price elasticity of demand of the consumers in the first group is equal to e1 and the price elasticity of demand for the consumers in the second group is equal to e2 and if the marginal cost of producing the good is constant and equal to "C" then the price charged by this monopoly firm to each one of the two groups is calculated by using the following formula

"\frac{P1}{P2}=\frac{1+\frac{1}{e2}}{1+\frac{1}{e1}}"

In the above equation "P1" represents the price charged by the monopoly firm to group 1 customers and "P2" represents the price charged by the monopoly firm to group 2 customers.